There are 850 Douglas fir and ponderosa pine trees in a
section of forest bought by Sawz Logging Co. The company paid an average of
$300 for each Douglas fir and $225 for each ponderosa pine. If the company paid
$217,500 for the trees, how many of each kind did the company buy?

Answer by tutorcecilia(2152)

Let x = Douglas Firs
Let y = Ponderosa Pines
Total trees = 850
So x+y=850 and
y=850-x
.
$300(x)+$225(y)=$217,500 (total cost of trees)
300x+225(850-x)=217,500 [substitute ((850-x) for the y-term and simplify]
300x+191,250-225x=217,500
75x+191,250=217,500 [solve for the x-term]
75x=217,000-191,2500
75x=26,250
x = 350 Douglas Firs
So, x+y=850
350+y=850
y=850-350=500 Ponderosa Pines
.
check by plugging the values for x and y back into the original equation and solve:
$300(350x)+$225(500)=$217,500
217,500=217,500 [checks out]